# Semilinear spaces of matrices with transitivity, rank or spectral constraints

### v1

@@ 1.2 Reduction to the irreducible case
Let $\mathcal{S}$ be an optimal $\mathbb{F}$-trivial spectrum (linear) subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$, with invariant subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$. Denote, for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, by $\mathcal{S}_{i}$ the set of all endomorphisms of $V_{i}/V_{i-1}$ induced by the elements of $\mathcal{S}$. Then
$\mathcal{S}=\mathcal{S}_{1}\vee\cdots\vee\mathcal{S}_{p},$
and for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ the set $\mathcal{S}_{i}$ is an irreducible optimal $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$.
Let us now give a matrix interpretation of the above. Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ are called similar, and we write $\mathcal{X}\simeq\mathcal{X}^{\prime}$, when there exists $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{D})$ such that $\mathcal{X}^{\prime}=P\mathcal{X}P^{-1}$.
Let $n_{1},\dots,n_{p}$ be positive integers. For respective subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, we denote by $\mathcal{X}_{1}\vee\cdots\vee\mathcal{X}_{p}$ the set of all matrices of the form
$M=\begin{bmatrix}M_{1}&&[?]\\ &\ddots&\\ [0]&&M_{p}\end{bmatrix}\quad\text{where $M_{1}\in\mathcal{X}_{1},\dots,M_{p}\in\mathcal{X}_{p}$.}$
We call it the joint of the sets $\mathcal{X}_{1},\dots,\mathcal{X}_{p}$.
Let $\mathcal{M}$ be an optimal $\mathbb{F}$-linear subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$. Then there exists a unique partition $n=n_{1}+\cdots+n_{p}$ into positive integers, and, for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, an irreducible optimal $\mathbb{F}$-linear subspace $\mathcal{M}_{i}$ of $\operatorname{\operatorname{M}}_{n_{i}}(\mathbb{D})$ such that
$\mathcal{M}\simeq\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}.$
Moreover, each $\mathcal{M}_{i}$ is uniquely determined by $\mathcal{M}$ up to similarity.
As a conclusion, we find that understanding the optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ essentially amounts to understanding the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$ for all integers $k\in\mathopen{[\![}1,n\mathclose{]\!]}$. There is a small difficulty however: given a partition $n=n_{1}+\cdots+n_{p}$ and respective irreducible optimal $\mathbb{F}$-trivial spectrum subspaces $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$ of $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, it is possible in theory that $\mathcal{M}_{1}\vee\dots\vee\mathcal{M}_{p}$, which is an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$, is not an optimal one. We call this the composition problem. Fortunately, in the situations we will consider, it will turn out that the joint of irreducible optimal $\mathbb{F}$-trivial spectrum subspaces is always optimal, and this will stem from the observation that
$\sum_{k=1}^{p}\alpha(n_{k},d)+d\sum_{1\leq i<j\leq p}n_{i}n_{j}=\alpha(n,d).$
1.3 Basic considerations on trivial spectrum subspaces, and a review of known results
Now, we recall the main known results on trivial spectrum subspaces in the case $\mathbb{D}=\mathbb{F}$.
Theorem 1.3 (See [12, 15]).
Let $V$ be a vector space over $\mathbb{F}$ with finite dimension $n$. Then the greatest possible dimension for an $\mathbb{F}$-trivial spectrum subspace of $\operatorname{\operatorname{End}}(V)$ is $\dbinom{n}{2}=\alpha(n,1)$.
In particular, the composition problem stated in the previous section has a positive answer when $\mathbb{D}=\mathbb{F}$.
As far as the irreducible optimal $\mathbb{F}$-trivial spectrum subspaces are concerned, their structure is described by the main theorem of [18] when $\mathbb{D}=\mathbb{F}$, which excludes the case $|\mathbb{F}|=2$ (in that case there are enough counterexamples to discourage anyone to try formulating a plausible conjecture [22]). Before we restate this theorem, some additional definitions and notation are required. Beware that throughout we manipulate bilinear forms that are possibly nonsymmetric.
Definition 1.1.
Let $b$ be a bilinear form on an $\mathbb{F}$-vector space $V$. We say that $b$ is nonisotropic whenever
$\forall x\in V,\;b(x,x)\neq 0.$
An endomorphism $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ is called $b$-alternating (on the right) whenever
$\forall x\in V,\;b(x,u(x))=0,$
i.e., the bilinear form $(x,y)\in V^{2}\mapsto b(x,u(y))$ is alternating.
We denote by $\mathcal{A}_{b}$ the set of all $b$-alternating $\mathbb{F}$-endomorphisms of $V$ (it is a linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$). If $V$ has an additional structure of $\mathbb{D}$-vector space (which extends its structure of $\mathbb{F}$-vector space), we set
$\mathcal{A}_{b,\mathbb{D}}:=\mathcal{A}_{b}\cap\operatorname{\operatorname{End}}_{\mathbb{D}}(V).$
If $V$ has dimension $n$ over $\mathbb{D}$ and $b$ is nondegenerate, then $\mathcal{A}_{b}$ is naturally isomorphic to the space of all alternating bilinear forms on $V$, and hence has dimension $\frac{nd(nd-1)}{2}$ over $\mathbb{F}$. Moreover, if $b$ is nonisotropic then the linear subspace $\mathcal{A}_{b}$ has $\mathbb{F}$-trivial spectrum, as for all $x\in V$ and $u\in\mathcal{A}_{b}$ the equality $u(x)=x$ implies $b(x,x)=b(x,u(x))=0$ and hence $x=0$. Now, we can state the main result of [18], reformulated in geometric terms.
Theorem 1.4 (See theorems 3 and 4 in [18]).
Let $V$ be a nonzero (finite-dimensional) $\mathbb{F}$-vector space, with $|\mathbb{F}|>2$. Then:
@@ 1.3 Basic considerations on trivial spectrum subspaces, and a review of known results
    The irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are the spaces of the form $\mathcal{A}_{b}$, where $b$ is a nonisotropic bilinear form on $V$.
@@
    Given two nonisotropic bilinear forms $b$ and $b^{\prime}$ on $V$, the equality $\mathcal{A}_{b}=\mathcal{A}_{b^{\prime}}$ holds if and only if $b^{\prime}$ is a scalar multiple of $b$, and $\mathcal{A}_{b}$ is similar [^3] Two subsets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are called similar when there exists a vector space automorphism $\varphi$ of $V$ such that $\mathcal{X}^{\prime}=\{\varphi\circ u\circ\varphi^{-1}\mid u\in\mathcal{X}\}$. to $\mathcal{A}_{b^{\prime}}$ if and only if $b^{\prime}$ is congruent to $b$ (i.e., equivalent to a scalar multiple of $b$).
Denote by $\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ the space of all alternating matrices of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ (a matrix $M$ is alternating if and only if it represents an alternating bilinear form in a basis, or in other words $M^{T}=-M$ and all the diagonal entries of $M$ are equal). In matrix terms, point (a) of Theorem 1.4 means that the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ are the spaces of the form $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ for the nonisotropic matrices $P\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$. And the second part of point (b) means that two such spaces $P^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ and $Q^{-1}\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ are similar if and only if $P$ is congruent to a scalar multiple of $Q$ (i.e., $P=\lambda\,RQR^{T}$ for some $\lambda\in\mathbb{F}^{\times}$ and some $R\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$). Moreover, for the more general case of (potentially reducible) optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$, we have the following corollary:
@@
Let $V$ be an $\mathbb{F}$-vector space, where $|\mathbb{F}|>2$, and $\mathcal{S}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$. Then the $\mathcal{S}$-invariant linear subspaces form a flag of subspaces $V_{0}\subset V_{1}\subset\cdots\subset V_{p}=V$, and there exists a list $(b_{1},\dots,b_{p})$ of nonisotropic bilinear forms, with $b_{i}$ defined on $(V_{i}/V_{i-1})^{2}$, such that $\mathcal{S}$ is the space of all $u\in\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ that leave each subspace $V_{i}$ invariant and induce a $b_{i}$-alternating endomorphism $u_{i}$ of $V_{i}/V_{i-1}$ for all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$. Moreover each form $b_{i}$ is uniquely determined up to multiplication by a nonzero scalar.
@@
In matrix terms this yields that for every optimal trivial spectrum subspace $\mathcal{M}$ of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$, there is a partition $n=n_{1}+\cdots+n_{p}$ into positive integers and a list $(P_{1},\dots,P_{p})\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F})\times\cdots\times\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ of nonisotropic matrices such that $\mathcal{M}$ is conjugated to the space of all matrices of the form
$\begin{bmatrix}P_{1}^{-1}A_{1}&[?]&&[?]\\ [0]&P_{2}^{-1}A_{2}&\ddots&\\ \vdots&\ddots&\ddots&[?]\ \\ [0]&\cdots&[0]&P_{p}^{-1}A_{p}\end{bmatrix}$
where $A_{1}\in\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}),\dots,A_{p}\in\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F})$ are variable alternating matrices, and the question marks represent arbitrary blocks. Moreover, the list $(n_{1},\dots,n_{p})$ is uniquely determined by $\mathcal{M}$, and each $P_{i}$ is uniquely determined by $\mathcal{M}$ up to congruence and multiplication with a nonzero scalar. Conversely, every space of the above form is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
Hence the structure of the optimal trivial spectrum subspaces is intimately connected with the quadratic structure of $\mathbb{F}$. If $\mathbb{F}$ is quadratically closed, irreducible optimal subspaces only exist in dimension $1$, and we obtain the same result as in Gerstenhaber’s theorem (simultaneous triangularization). If $\mathbb{F}$ is finite then irreducible optimal subspaces only exist in dimensions $1$ and $2$, and there is a very limited number of similarity classes of such spaces. If $\mathbb{F}$ is the field of reals, there are irreducible optimal subspaces in all dimensions, and their classification up to conjugation is rather easy (it can be reduced to the determination of the orbits, under orthogonal conjugation, of the skewsymmetric matrices).
1.4 Main new results on trivial spectrum subspaces
We will now state a generalization of the previous results, however with a cardinality assumption on $\mathbb{F}$.

### v2

@@ 1.2 Reduction to the irreducible spaces
Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V)$ for some vector space $V$. Then:
@@
    The $S$-invariant subspaces of $V$ are totally ordered under inclusion, thereby forming a (potentially incomplete) flag $(V_{0},\dots,V_{p})$ of linear subspaces of $V$.
@@
    For all $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, the $\mathbb{F}$-linear subspace $S_{i}$ of $\operatorname{\operatorname{End}}_{\mathbb{D}}(V_{i}/V_{i-1})$ induced by $S$ is an irreducible optimal trivial spectrum subspace.
  3. (c)
    The space $S$ equals $S_{1}\vee\cdots\vee S_{p}$.
In matrix terms, this translates as follows:
@@
Let $\mathcal{M}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ for some integer $n>0$. Then there is a partition $n=n_{1}+\cdots+n_{p}$ into positive integers, and a list $(\mathcal{M}_{1},\dots,\mathcal{M}_{p})$ in which:
  1. (i)
    For each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$, the set $\mathcal{M}_{i}$ is an irreducible optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n_{i}}(\mathbb{D})$.
  2. (ii)
    The space $\mathcal{M}$ is similar to $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$.
  3. (iii)
    The similarity class of each $\mathcal{M}_{i}$ is uniquely determined by the one of $\mathcal{M}$.
This shows that the problem of classifying, up to conjugation, the optimal trivial spectrum subspaces of matrices over $\mathbb{D}$ can be reduced to the subproblem of classifying the irreducible ones.
One difficulty must be pointed out though, and we call it the reconstruction problem. If we have irreducible optimal trivial spectrum subspaces of matrices $\mathcal{M}_{1},\dots,\mathcal{M}_{p}$, then although their joint $\mathcal{M}_{1}\vee\dots\vee\mathcal{M}_{p}$ is surely a trivial spectrum subspace, there can be no simple way to prove that it is optimal. To see this, consider the mundane case where we take $\mathbb{F}$-linear hyperplanes $H_{1},\dots,H_{n}$ of $\mathbb{D}$ that do not contain $1_{\mathbb{D}}$, and we form the joint $H_{1}\vee\cdots\vee H_{n}$. Clearly each $H_{i}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{1}(\mathbb{D})$, and their joint can be optimal only if we prove that the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{D})$ is $\alpha(n,d)$. Yet, as we shall see this result is not known when $\mathbb{F}$ is an arbitrary finite field (unless $d=1$).
However, if we have an integer $n>0$ and we know that $\alpha(k,d)$ is the greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{k}(\mathbb{D})$ for all $k\in\mathopen{[\![}1,n\mathclose{]\!]}$ then for any partition $n=n_{1}+\cdots+n_{p}$ and any list $(\mathcal{M}_{1},\dots,\mathcal{M}_{p})$ of optimal trivial spectrum subspaces of, respectively, $\operatorname{\operatorname{M}}_{n_{1}}(\mathbb{D}),\dots,\operatorname{\operatorname{M}}_{n_{p}}(\mathbb{D})$, the space $\mathcal{M}_{1}\vee\cdots\vee\mathcal{M}_{p}$ is optimal since it is easily observed that
$n(d-1)+d\dbinom{n}{2}=\sum_{i=1}^{p}\left(n_{i}(d-1)+d\dbinom{n_{i}}{2}\right)+\sum_{1\leq i<j\leq p}dn_{i}n_{j}.$
1.3 Review of older results
Now that the main concepts are in place, we can review the known results on our problem. As said in the introduction, these results deal with the special case $\mathbb{D}=\mathbb{F}$ only:
Theorem 1.3 (Quinlan [9], de Seguins Pazzis [10]).
Let $n>0$. The greatest possible dimension for a trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ is $\dbinom{n}{2}$.
The structure of the spaces of greatest possible dimension is connected with bilinear forms. Let $V$ be an $\mathbb{F}$-vector space and $b$ be a bilinear form on $V$. An endomorphism $u$ of $V$ is called $b$-alternating (on the right) whenever $\forall x\in V,\;b(x,u(x))=0$, i.e. the bilinear form $(x,y)\in V^{2}\mapsto b(x,u(y))$ is alternating.
Notation 1.2.
We denote by $\mathcal{A}_{b}$ the set consisting of all the $b$-alternating endomorphisms of $V$. This is a linear subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$.
Assume now that $b$ is nonisotropic, that is $\forall x\in V,\;b(x,x)\neq 0$. Then $\mathcal{A}_{b}$ has dimension $\frac{n(n-1)}{2}$ (by the non-degeneracy of $b$), and it is a trivial spectrum subspace (indeed, for all $u\in\mathcal{A}_{b}$, having $u(x)=x$ for some $x\in V\smallsetminus\{0\}$ implies that $0=b(x,u(x))=b(x,x)$). It is not difficult to prove that it is also an irreducible one if $n>0$.
Theorem 1.4 (de Seguins Pazzis [12]).
Assume that $|\mathbb{F}|>2$. Let $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n>0$. Then the irreducible optimal trivial spectrum subspaces of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ are the spaces of the form $\mathcal{A}_{b}$ for some nonisotropic bilinear form $b$ on $V$.
For two such forms $b$ and $b^{\prime}$, the spaces $\mathcal{A}_{b}$ and $\mathcal{A}_{b^{\prime}}$ are equal if and only if $b^{\prime}=\lambda b$ for some $\lambda\in\mathbb{F}^{\times}$, and they are conjugated in $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ if and only if $b^{\prime}$ is congruent to $b$ (i.e., there exists an automorphism $\varphi$ of $V$ and a scalar $\alpha\in\mathbb{F}^{\times}$ such that $b^{\prime}(x,y)=\alpha\,b(\varphi(x),\varphi(y))$ for all $x,y$ in $V$).
As a consequence, we get the full picture for optimal trivial spectrum subspaces:
@@ 1.3 Review of older results
Assume that $|\mathbb{F}|>2$. Let $V$ be an $n$-dimensional vector space over $\mathbb{F}$, with $n>0$. Let $S$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$, with flag of invariant subspaces denoted by $(V_{0},\dots,V_{p})$. Then for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ there is a nonisotropic bilinear form $b_{i}$ on $V_{i}/V_{i-1}$, uniquely determined by $S$ up to multiplication with a nonzero scalar, such that $S=\mathcal{A}_{b_{1}}\vee\cdots\vee\mathcal{A}_{b_{p}}$.
Conversely, given a flag $(V_{0},\dots,V_{p})$ of linear subspaces of $V$, and for each $i\in\mathopen{[\![}1,p\mathclose{]\!]}$ a nonisotropic bilinear form $b_{i}$ on $V_{i}/V_{i-1}$, the set $S:=\mathcal{A}_{b_{1}}\vee\cdots\vee\mathcal{A}_{b_{p}}$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{End}}_{\mathbb{F}}(V)$ with associated flag $(V_{0},\dots,V_{p})$ of invariant subspaces.
We can also state the matrix version of the problem, where a square matrix $P\in\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ is called nonisotropic whenever $\forall X\in\mathbb{F}^{n}\smallsetminus\{0\},\;X^{T}PX\neq 0$, and two square matrices $M$ and $M^{\prime}$ are called homocongruent whenever there exist a scalar $\alpha\in\mathbb{F}^{\times}$ and an invertible matrix $Q\in\operatorname{\operatorname{GL}}_{n}(\mathbb{F})$ such that $M^{\prime}=\alpha\,QMQ^{T}$. Also, for $n\in\mathbb{N}^{*}$ we denote by $\operatorname{\operatorname{A}}_{n}(\mathbb{F})$ the space of all alternating $n$-by-$n$ matrices (i.e., skew-symmetric matrices with all diagonal entries zero).
Theorem 1.6.
Assume that $|\mathbb{F}|>2$.
  1. (a)
    Let $n>0$ and $\mathcal{M}$ be an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Then there exist a partition $n=n_{1}+\cdots+n_{p}$ and nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ such that M(P_1^-1 A_n_1(F)) ∨⋯∨(P_p^-1 A_n_p(F)). The partition $(n_{1},\dots,n_{p})$ is then uniquely determined by $n$, as well as the homocongruence classes of the matrices $P_{1},\dots,P_{p}$.
  2. (b)
    Conversely, for each partition $n=n_{1}+\cdots+n_{p}$ and each list of nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$, the set $(P_{1}^{-1}\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}))\vee\cdots\vee(P_{p}^{-1}\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F}))$ is an optimal trivial spectrum subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
Finally, we can state the solution of the initial problem of classifying the optimal unital affine subspaces of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Here, the matrix formulation turns out to be the clearer one:
Theorem 1.7.
Assume that $|\mathbb{F}|>2$.
  1. (a)
    Let $n>0$ and $\mathcal{M}$ be an optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$. Then there exist a partition $n=n_{1}+\cdots+n_{p}$ and nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$ such that M(P_1+ A_n_1(F)) ∨⋯∨(P_p+ A_n_p(F)). The partition $(n_{1},\dots,n_{p})$ is then uniquely determined by $n$, as well as the congruence class of each quadratic form $X\mapsto X^{T}P_{i}X$.
  2. (b)
    Conversely for each partition $n=n_{1}+\cdots+n_{p}$ and each list of nonisotropic matrices $P_{1}\in\operatorname{\operatorname{GL}}_{n_{1}}(\mathbb{F}),\dots,P_{p}\in\operatorname{\operatorname{GL}}_{n_{p}}(\mathbb{F})$, the set $(P_{1}+\operatorname{\operatorname{A}}_{n_{1}}(\mathbb{F}))\vee\cdots\vee(P_{p}+\operatorname{\operatorname{A}}_{n_{p}}(\mathbb{F}))$ is an optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$.
Hence, this theorem gives a one-to-one correspondence between the equivalence classes of the optimal unital affine subspace of $\operatorname{\operatorname{M}}_{n}(\mathbb{F})$ on the one hand, and the lists of congruence classes of nonisotropic quadratic forms of respective dimensions $n_{1},\dots,n_{p}$ such that $n=\sum_{k=1}^{p}n_{k}$ on the other hand.
Remark 1.1.
In the previous classification results, the case of fields with two elements is systematically omitted. Indeed, it has been shown through various examples (see section 5.2 of [17]) that a classification theorem for $\mathbb{F}_{2}$ is hopeless.
1.4 The new results on trivial spectrum subspaces
We are now ready to state the results on the more general case of a simple $\mathbb{F}$-algebra. From now on, we consider a division ring $\mathbb{D}$ with finite dimension $d$ over a central subfield $\mathbb{F}$.
We start with the statement on the greatest possible dimension for a trivial spectrum (linear) subspace. This generalizes Theorem 1.3, however with a cardinality assumption on $\mathbb{F}$.
Theorem 1.8.